The management of JAL Airlines wants to determine if there is a linear relationship between advertising expenditures
and the number of passengers who choose to fly JAL.

The following data is collected over the past 15 months of performance by JAL Airlines.
Note that ADEXP represents Advertising Expenditures in $1,000s, and Passengers is the
number of passengers in 1,000s).

ADEXP

PASSENGERS

100

15

120

17

80

13

170

23

100

16

150

21

100

14

140

20

190

24

100

17

110

16

130

18

160

23

100

15

120

16

a)    Find the regression equation relating Passengers to ADEXP.

b)   Calculate the correlation coefficient, accurate to three decimal places  

c) Test the significance of the correlation coefficient with a 0.05 level of significance.
Include definition of hypotheses, test statistic, p-value or critical value, decision about Ho, and conclusion.

d)  What is the best estimate for number of passengers when $125,000 is spent on ads?
Answer #1

Suppose, random variables X and Y denote expenditure (in thousand dollars) in advertisements and number of passengers who choose to fly JAL Airlines respectively.

(a)

Regression equation of y on x is given by

\tiny y-\bar y=b_{yx}\left(x-\bar x\right)

Here,

Number of pairs of observation \tiny n=15

\tiny \bar y=\frac{1}{n}\sum_{i=1}^{n}y_i=\frac{1}{15}\sum_{i=1}^{15}y_i=17.86667

\tiny \bar x=\frac{1}{n}\sum_{i=1}^{n}x_i=\frac{1}{15}\sum_{i=1}^{15}x_i=124.6667

\tiny b_{yx}=\frac{Cov(X,Y)}{Var(X)}=0.1081317

\tiny \therefore y-17.86667=0.1081317\left(x-124.6667\right)

\tiny \Rightarrow y-17.86667=0.1081317x-0.1081317*124.6667

\tiny \Rightarrow y=0.1081317x+17.86667-13.48042

\tiny \therefore y=0.1081317x+4.38625

(b)

Correlation coefficient is given by

\tiny r_{x,y}=\frac{Cov(X,Y)}{\sqrt {Var(X)*Var(Y)}}=0.9683784

(c)

We have to test for null hypothesis \tiny H_0:\rho_{xy}=0

against the alternative hypothesis \tiny H_1:\rho_{xy}\neq0

Our test statistic is given by

\tiny t=\frac{r_{xy}\sqrt{n-2}}{\sqrt{1-r_{xy}^2}}

Here,

Sample size \tiny n=15

Sample correlation coefficient is given by

\tiny r_{x,y}=\frac{Cov(X,Y)}{\sqrt {Var(X)*Var(Y)}}=0.9683784

\tiny \therefore t_{calculated}=\frac{0.9683784\sqrt{13}}{\sqrt{1-0.9683784^2}}=13.99492

Degrees of freedom \tiny v=n-2=13

\tiny \text{p-value}=P\left(t_{13}\leq-13.99492\right)+P\left(t_{13}\geq13.99492\right)=0.000000003238011

[Using R-code 'pt(-13.99492,13)+1-pt(13.99492,13)']

Level of significance \tiny \alpha=0.05

We reject our null hypothesis if \tiny \text{p-value}<\alpha

Here, we observe that \tiny \text{p-value}=0.000000003238011\nless0.05=\alpha

So, we reject our null hypothesis.

Hence, based on the given data we can conclude that there is significant evidence that there is a linear relationship between advertising expenditures and the number of passengers who choose to fly JAL.

(d)

For \tiny x=125 , predicted number of passengers is given by

\tiny \hat y=0.1081317*125+4.38625=17.90271

Hence, best estimate for number of passengers is 18.